OpenStudy (anonymous):

Help Please: A culture started with 3000 bacteria. After 4 hours, it grew to 3,600 bacteria. Predict how many bacteria will be present after 10 hours. Round your answer to the nearest whole number. Formula to follow; P=Ae^kt

OpenStudy (anonymous):

So you will need to solve for k in this case and then plug it back in I don't have a calulator on me right now but try this.. 3000=3600e^k*4 solving for k (growth constant) would give you 3000/3600=k*4 ln(3000)/3600)/4=k this should give you your rate constant, now plug the k value in for the following => 3600=Ae^k*10 Now solve for A and this should give you your answer.. ithink?

OpenStudy (anonymous):

here try using this because I can figure out how to get is, this is a really great calculator: https://www.mathway.com/

OpenStudy (anonymous):

ok so 5,679 ithink?

OpenStudy (anonymous):

alright i just got locked out of my assignment but still here is a medal for helping (my connection timed out :( )

OpenStudy (anonymous):

gah, lemee know and I can explain it better for you.